Can someone help me on this problem?
I need to solve the following word problem. I need to also label the unknowns and to show the equation you use for the solution:
The Randolphs used 12 more gal. of fuel oil in October than in September and twice as much oil in November as in September. If they used 132 gal. for the 3 months, how much was used during each month?
Thanks, Sher

Answer by checkley71(8403)

SEPT=X
OCT=X+12
NOV=2*X
THUS X+(X+12)+2*X=132 OR X+X+12+2X=132 OR 4X+12=132 4X=120 OR X=30 GAL.
THEREFORE THE SEPT USE=30 GAL. THE OCT. USE=30+12=42 GAL. & NOV. USE=2*30=60 GAL.
PROOF 30+42+60=132 OR 132=132